OpenStudy (anonymous):

An angle is 3 more than twice the measure of its compliment. What is the measure of the larger angle?

OpenStudy (anonymous):

let the complement be x then the angle is 2x+3 x+2x+3=90 or 3x=90-3 3x=87 x=29 hence the angle is 90-29=61